# Minimal cellular automaton model with heterogeneous cell sizes predicts epithelial colony growth

### v1

@@ IV Discussion
We developed a mechanistic, data-driven model to simulate the dynamics of planar cell cultures by extending a probabilistic cellular automaton to incorporate size changes of individual cells during growth and cell division. We successfully apply this CA model to previous in-vitro experiments PulHufNevStrSigFygShr2012 of epithelial tissue composed of Madin-Darby canine kidney cells. In particular, the model parameters are calibrated using measurements based on single-cell tracking and subsequently the resulting CA model is validated against independent experiments on the development of culture size, cell density, and cell velocity. The agreement between model predictions and these experimental observations suggests that the exclusion principle as sole inter-cellular interaction paired with a size-dependent proliferation rate of each cell is sufficient to explain the emerging dynamics of contact inhibition. In the CA model, the reproduced transition from exponential to quadratic growth of the colony originates from the increasing lack of space in the center of the colony, inhibiting cell growth and division, when the expansion of the colony can no longer be compensated for by the migration of the exterior cells. This is consistent with previous interpretations and the observation that contact inhibition can take place several days after cells have been in contact PulHufNevStrSigFygShr2012 ; AlaHatLowVoi2015 . Moreover, the collective, outwards motion of the exterior cells and the rapidly increasing cell density in the center of the culture are a consequence of these simple mechanical constraints.
@@
As mentioned there is a wide variety of theoretical models to describe and simulate the development of healthy and cancerous cellular tissue and even contact inhibition GaoMcDHlaEnd2013 ; AndCha2012 ; SciPreWol2013 ; SciPre2012 ; SotVerTav2011 ; GivSciPreBuoFun2010 ; RejAnd2011 ; AhmGra2018 ; NobBurLeLemVioKatBee2022 ; PodDelElg2015 . Most of these models explicitly consider intra- and inter-cellular forces, allowing to study the delicate interplay between biomechanical quantities like tension, pressure, adhesion, as well as cellular stiffness and compressibility, which govern cell topology and large scale phenomena like elastic, plastic, viscoelastic deformations. For example, previously a mechanistic hybrid continuum-discrete model has been applied to the same experiment of epithelial cell colonies as our study AlaHatLowVoi2015 . Besides culture size, cell density and cell area median, this model was able to reproduce the observed cell topology and radial distribution function. However, the spatial scale of the culture had to be fitted empirically to the culture size, as the model underpredicts the cluster size where the transition to a stationary, postconfluent state occurs, and the resulting cell velocities are reported to be ten times smaller than in the experiment. This is partially due to computational cost, which prevented the use of a larger parameter of mobility, which also reflects effects of cell-substrate adhesion. This example highlights two more general issues of the above-mentioned types of models, see also Ref. Vos2012 : Firstly, they often rely on empirical or effective parameters, like coefficients of free energy functions or elastic properties of tissue, which are either biologically not known or at least very difficult to measure, in contrast to the properties of single cells. Secondly, they often entail high computational effort, which limits exploration of the parameter space and consequently the calibration to concrete experimental data.
@@
While CA models have the advantage, that their parameters are usually based on biologically motivated, single-cell behavior and that they are computationally relatively cheap, the incorporation of biomechanical forces has been a challenge with few exceptions like cell adhesion FraAlaBoeVosLan2022 ; RosBoeLanVos2021 . The type of CA we propose here has the potential to partially overcome this challenge: The fact that cells of different sizes can be located at the same grid point naturally defines local cell densities and density gradients. Together with the possibility to define cellular, size-dependent mechanisms and forces, beyond proliferation and growth, this allows to effectively introduce biomechanical processes from other types of models, e.g., homeostatic pressure PodDelElg2015 ; BasRisJoaSasPro2009 or cell bulk stress in terms of cell compression AlaHatLowVoi2015 from the ratio of the actual size of a cell and its (dynamic) target size. Furthermore, in our CA model cells are capable of moving past each other, which is typically not inherent to both common CA models and vertex models. This feature can be exploited to study effects like collective motion or fingering SzaMer2013 ; SzaVarGarHegCzi2012 by adding direction vectors of persistent motion to each cell, which are affected by neighboring cell motion or cell densities. Moreover, the concept of local cell densities allows to couple the CA model to additional PDEs to incorporate the interaction of cells with external fields like oxygen, nutrients, and metabolic waste. Finally, the CA model can be directly extended to three-dimensional space. While none of these concepts were in the end necessary for the application of the CA model to the considered experiments of epithelial growth, they may guide efficient, data-driven modeling of future scenarios of tissue dynamics.
V Acknowledgement
@@
We are grateful for discussions with S. Rühle, M. Luft and F. Freier. The authors acknowledge that this research has been co-financed by the EU, the European Social Fund (ESF), and by tax funds on the basis of the budget passed by the Saxon state parliament (project SAB-Nr. 100382145) and the Bundesministerium für Bildung und Forschung (BMBF 16dkwn001a/b). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
VI Conflicts of Interest
@@
VII Data Availability
All relevant data are within the manuscript and its Supporting information files.
@@ Appendix A Supplemental figures
Figure 11: CA model displays transition from undirected, uniform, fast cell motion to outwards biased motion with radial velocity gradient (from top to bottom, note shrinking scale bar). Panels analogous to Fig. 10 in the main text, but for different times around the transition point $t_{T}=6\pm 1$ d.
@@
Figure 12: Emergent dynamics of calibrated CA model (red line) reproduces experimentally observed size of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 1B)). The colony size exhibits initially an exponential growth (low-density, confluent regime) and slows around time $t_{T}=6\pm 1$ d to a subexponential growth (transition to a stationary, postconfluent state). Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region. Model parameters as Fig. 6 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}^{2}/\text{h}$.
@@
Figure 13: Emergent dynamics of calibrated CA model (red line) reproduces experimentally observed cell densities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 1C), corresponding data to areas displayed in Fig. 12). The cell density is initially constant (low-density, confluent regime) and increases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state) implying dense packing of cells, which become smaller and smaller due to slowing but continued proliferation without cell growth. Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region. Model parameters as Fig. 7 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}^{2}/\text{h}$.
@@
Figure 14: Emergent dynamics of calibrated CA model (red line) reproduces qualitatively the experimentally observed cell velocities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 2C), corresponding data to Figs. 12 and 13). The average cell velocity is initially constant (low-density, confluent regime) and decreases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state). Presented are the root-mean-square velocities, where for the model velocities are evaluated from a single simulation run over time increments of $580$ min to neglect small-scale oscillations and standard deviation is displayed as red shaded region. Model parameters as Fig. 9 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}^{2}/\text{h}$.
@@
Figure 15: Emergent dynamics of calibrated CA model (red line) consistent with experimentally observed cell size median for initially uniformly seeded cells (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 3C), transition $t_{T}$ to a stationary, postconfluent state corresponds roughly to $1.5$ d). Cell size median initially drops rapidly and then saturates at small cell sizes $A\sim 40\ \mu\text{m}^{2}$. Model parameters as Fig. 8 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}^{2}/\text{h}$. Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region.

### v2

@@ V Discussion
We developed a mechanistic, data-driven model to simulate the dynamics of planar cell cultures by extending a probabilistic cellular automaton to incorporate size changes of individual cells during growth and cell division. We successfully apply this CA model to previous in-vitro experiments PulHufNevStrSigFygShr2012 of epithelial tissue composed of Madin-Darby canine kidney cells. In particular, the model parameters are calibrated using measurements based on single-cell tracking and subsequently the resulting CA model is validated against independent experiments on the development of culture size, cell density, and cell velocity. The agreement between model predictions and these experimental observations suggests that the exclusion principle as sole inter-cellular interaction paired with a size-dependent proliferation rate of each cell is sufficient to explain the emerging dynamics of contact inhibition. In particular, incorporation of additional mechanisms like cell-cell signaling or cell adhesion are not necessary. In the CA model, the reproduced transition from exponential to quadratic growth of the colony originates from the increasing lack of space in the center of the colony, inhibiting cell growth and division, when the expansion of the colony can no longer be compensated for by the migration of the exterior cells. This origin of the transition is consistent with previous interpretations and the observation that contact inhibition can take place several days after cells have been in contact PulHufNevStrSigFygShr2012 ; AlaHatLowVoi2015 . Moreover, the collective, outwards motion of the exterior cells and the rapidly increasing cell density in the center of the culture are a consequence of these simple mechanical constraints.
@@
As mentioned there is a wide variety of theoretical models to describe and simulate the development of healthy and cancerous cellular tissue and even contact inhibition GaoMcDHlaEnd2013 ; AndCha2012 ; SciPreWol2013 ; SciPre2012 ; SotVerTav2011 ; GivSciPreBuoFun2010 ; RejAnd2011 ; AhmGra2018 ; NobBurLeLemVioKatBee2022 ; PodDelElg2015 . Most of these models explicitly consider intra- and inter-cellular forces, allowing to study the delicate interplay between biomechanical quantities like tension, pressure, adhesion, as well as cellular stiffness and compressibility, which govern cell topology, e.g., the distribution of the number of direct cell neighbors, and large scale phenomena like elastic, plastic, viscoelastic deformations. For example, previously a mechanistic hybrid continuum-discrete model has been applied to the same experiment of epithelial cell colonies as our study AlaHatLowVoi2015 . Besides culture size, cell density and cell area median, this model was able to reproduce the observed cell topology and radial distribution function. However, the spatial scale of the culture had to be fitted empirically to the culture size, as the model underpredicts the cluster size where the transition to a stationary, postconfluent state occurs, and the resulting cell velocities are reported to be ten times smaller than in the experiment. This is partially due to computational cost, which prevented the use of a larger parameter of mobility, which also reflects effects of cell-substrate adhesion. This example highlights two more general issues of the above-mentioned types of models, see also Ref. Vos2012 : Firstly, they often rely on empirical or effective parameters, like coefficients of free energy functions or elastic properties of tissue, which are either biologically not known or at least very difficult to measure, in contrast to the properties of single cells. Secondly, they often entail high computational effort, which limits exploration of the parameter space and consequently the calibration to concrete experimental data.
@@
While CA models have the advantage, that their parameters are usually based on biologically motivated, single-cell behavior and that they are computationally relatively cheap, the incorporation of biomechanical forces has been a challenge with few exceptions like cell adhesion FraAlaBoeVosLan2022 ; RosBoeLanVos2021 . The type of CA we propose here has the potential to partially overcome this challenge: The fact that cells of different sizes can be located at the same grid point naturally defines local cell densities and density gradients. Together with the possibility to define cellular, size-dependent mechanisms and forces, beyond proliferation and growth, this allows to effectively introduce biomechanical processes from other types of models, e.g., homeostatic pressure PodDelElg2015 ; BasRisJoaSasPro2009 or cell bulk stress in terms of cell compression AlaHatLowVoi2015 from the ratio of the actual size of a cell and its (dynamic) target size. Note that the concept of local densities has already been exploited in previous spatial stochastic models, which allow multiple particles in each lattice site BarMan1996 ; SmiGri2016 , e.g., to define migration probabilities based on the density differences between lattice sites, which is similar to our CA model if cell size and cell growth is omitted. Mean-field approximations of these models lead to partial differential equations of the reaction-advection-diffusion type exhibiting matching dynamics SmiGri2016 ; CiaSmiGri2017 . However, these models neglect cell growth, a crucial effect which complicates the derivation of a corresponding continuum model. An example is the already mentioned hybrid continuum-discrete model, derived from a stochastic particle model using the framework of dynamic density functional theory AlaHatLowVoi2015 .
Furthermore, in our CA model cells are capable of moving past each other, which is typically not inherent to both common CA models and vertex models. This feature can be exploited to study effects like collective motion or fingering SzaMer2013 ; SzaVarGarHegCzi2012 by adding direction vectors of persistent motion to each cell, which are affected by neighboring cell motion or cell densities. This could help to reproduce the finger-like protrusions observed for the expanding epithelial colony PulHufNevStrSigFygShr2012 , which neither our CA model nor the corresponding vertex model AlaHatLowVoi2015 exhibit. Moreover, the concept of local cell densities allows to couple the CA model to additional PDEs to incorporate the interaction of cells with external fields like oxygen, nutrients, and metabolic waste. Finally, the CA model can be directly extended to three-dimensional space. However, features of small spatial scales like cell shape AlaHatLowVoi2015 and correlation lengths AlaHatLowVoi2015 ; PulHufNevStrSigFygShr2012 can not be captured by our model. While none of these concepts were in the end necessary for the application of the CA model to the considered experiments of epithelial growth, they may guide efficient, data-driven modeling of future scenarios of tissue dynamics.
VI Acknowledgement
@@
We are grateful for discussions with S. Rühle, M. Luft and F. Freier. The authors acknowledge that this research has been co-financed by the EU, the European Social Fund (ESF), and by tax funds on the basis of the budget passed by the Saxon state parliament (project SAB-Nr. 100382145), the Bundesministerium für Bildung und Forschung (BMBF 16dkwn001a/b), and the Sächsisches Staatsministerium für Wissenschaft und Kunst (SMWK) (project FORZUG II TP 3). The publication of this article is funded by the Open Access Publication Fund of Hochschule für Technik und Wirtschaft Dresden - University of Applied Sciences. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
VII Conflicts of Interest
@@
VIII Data Availability
All relevant data are within the manuscript and its Supporting information files. The code is publicly available under https://github.com/j-schmied/cellularautomata.
@@ Appendix A Supplemental figures
Figure 1: CA model displays transition from undirected, uniform, fast cell motion to outwards biased motion with radial velocity gradient (from top to bottom, note shrinking scale bar). Panels analogous to Fig. 10 in the main text, but for different times around the transition point $t_{T}=6\pm 1$ d. For a radial profile of the velocities see Fig. 2.
Figure 2: Average cell velocities in the calibrated CA with respect to distance $r$ from the center of the colony display radial velocity gradient (top panel) and radially outwards biased cell motion (middle and bottom panel). Shown are averaged the (top panel) absolute velocities $v\right>$, (middle panel) relative radial velocity $\frac{\vec{v}\cdot\vec{r}}{v\cdot r}\right>$, and (bottom panel) relative angular velocity $\frac{|\vec{v}\times\vec{r}|}{v\cdot r}\right>$ at the time points $4,5,6,7,10$ d corresponding to Fig. 1 and Fig. 10. While the angular velocity is always around zero (indicated by the dashed black line in the bottom panel), the radial velocity exhibits values well above zero.
Figure 3: Emergent dynamics of the calibrated CA model (red triangles) reproduce the observed division times $\tau_{2}$ for an initially homogeneously seeded cells (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 4E), blue line corresponds to the fitted proliferation rate from Fig. 5). The division times of the CA model are averaged over intervals of the cell size $A$ and corresponding standard deviation, resulting from the underlying exponential distribution of the division times, is shown as red shaded region. In the inset the according distributions of the division times of CA model (red) and experiment (black  (PulHufNevStrSigFygShr2012, , Fig. 4E inset)) are compared. While division times in the CA model follow an exponential distribution, biological cells divide at more regular intervals leading typically to a narrower distribution.
@@
Figure 4: Emergent dynamics of calibrated CA model (red line) reproduces experimentally observed size of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 1B)). The colony size exhibits initially an exponential growth (low-density, confluent regime) and slows around time $t_{T}=6\pm 1$ d to a subexponential growth (transition to a stationary, postconfluent state). Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region. Model parameters as Fig. 6 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}/\text{h}$.
@@
Figure 5: Emergent dynamics of calibrated CA model (red line) reproduces experimentally observed cell densities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 1C), corresponding data to areas displayed in Fig. 4). The cell density is initially constant (low-density, confluent regime) and increases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state) implying dense packing of cells, which become smaller and smaller due to slowing but continued proliferation without cell growth. Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region. Model parameters as Fig. 7 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}/\text{h}$.
@@
Figure 6: Emergent dynamics of calibrated CA model (red line) reproduces qualitatively the experimentally observed cell velocities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 2C), corresponding data to Figs. 4 and 5). The average cell velocity is initially constant (low-density, confluent regime) and decreases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state). Presented are the root-mean-square velocities, where for the model velocities are evaluated from a single simulation run over time increments of $580$ min to neglect small-scale oscillations and standard deviation is displayed as red shaded region. Model parameters as Fig. 9 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}/\text{h}$.
@@
Figure 7: Emergent dynamics of calibrated CA model (red line) consistent with experimentally observed cell size median for initially homogeneously seeded cells (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 3C), transition $t_{T}$ to a stationary, postconfluent state corresponds roughly to $1.5$ d). Cell size median initially drops rapidly and then saturates at small cell sizes $A\sim 40\ \mu\text{m}^{2}$. Model parameters as Fig. 8 in the main text, but for a smaller maximal migration velocity $v_{\text{max}}=50\ \mu\text{m}/\text{h}$. Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region.